OpenStudy (anonymous):

Can someone help me understand this? 2(-3a^8b)^3

OpenStudy (anonymous):

think of the stuff inside the bracket as x.So the question is 2x^3

OpenStudy (anonymous):

now exponentiation has higher priority tham multiplication. so first you gotta do x^3

OpenStudy (campbell_st):

ok... so the power of 3 operates on everything inside the brackets the power of a power law is \[(x^a)^b = x^{a \times b}\] you need to apply this law as well as \[2 \times (-3)^3\]

OpenStudy (anonymous):

I ended up getting -6^(3)2a^(11)2b^(3)

OpenStudy (anonymous):

wait up..

OpenStudy (anonymous):

first tell me what you get by doing x^3

OpenStudy (anonymous):

I am pretty confused right now..

OpenStudy (anonymous):

:D.ok i'll help you out

OpenStudy (campbell_st):

not quite.... the expanded form is \[2 \times (-3)^3 \times (a^8)^3 \times (b^1)^3\] apply the index law for power of a power, shown above... and evaluate the rest

OpenStudy (anonymous):

@nkswizzle did you get it mate?

OpenStudy (anonymous):

@dlearner struggling.. if you can solve it by step by step maybe I can get it :[.

OpenStudy (anonymous):

right.. (-3)^3=-27 (a^8)^3 = a^24 b^3 =b^3 :) now write it all together and multiply by tw

OpenStudy (anonymous):

@nkswizzle what did you get :)

OpenStudy (anonymous):

@dlearner I got 2 (-27)(a^24)(b^3) !

OpenStudy (anonymous):

put the two withem mate :)

OpenStudy (anonymous):

@dlearner Now I got -54 2a^24 2b^3

OpenStudy (anonymous):

it was simple was't it? :)

OpenStudy (anonymous):

hey yo>>> multiply the 2 only once @nkswizzle . its common to all

OpenStudy (anonymous):

@dlearner ah thanks ! It was just a bit confusing with the 2a^24 and the 2b^3 :p